Aaron is painting three rooms in his house. He has narrowed his paint choices down to eight. If he can choose the same color for more than one room, how many paint combinations does Aaron have?
Found 2 solutions by Finavon, sundararajchinnasamy:
Answer by Finavon(81)

In any room room he has an independent choice of 8 colours.
Choices = 8x8x8 = 2^9 = 512

Answer by sundararajchinnasamy(2)

given hint more than one room
therefore may be two of three rooms having same colour,or three rooms having same colour.
for two room having same color
3C2 =3 ways to select two rooms for having same color.now color arrangement is 3C2*8P2=168
for three rooms having same color
3C3*8P1=8
ans is 168+8=176